# Fast and accurate analytical formulas for light propagation in general static, spherically symmetric spacetimes

### v1

@@ I Introduction
Among these approaches, the Beloborodov formula (Eq. (1) in Beloborodov (2002)) stands out. In the Schwarzschild spacetime, it provides a relation between the emission angle of a light ray—measured with respect to the radial direction—and the angular (and radial) position at which the same point is observed in the asymptotic region. Despite its success, this formula has some intrinsic limitations: it loses accuracy for emission angles exceeding $90^{\circ}$, it assumes exact spherical symmetry (thus neglecting spin-induced spacetime deformations), and it is restricted to the Schwarzschild metric.
Some of these shortcomings have been mitigated in later works. In particular, Poutanen Poutanen (2020) proposed a refined expression—still within the Schwarzschild geometry—that significantly improves the treatment of large-angle emission and has been successfully applied in various astrophysical contexts.
Moreover, comparative analyses between approximate methods and full Kerr ray-tracing simulations have shown that, for moderate spins, both the resulting direct images and polarization signatures remain in close agreement Cárdenas-Avendaño et al. (2023); Loktev, Vladislav et al. (2024). Notably, the Poutanen approximation has been implemented in the ARTPOL algorithm, which enables fast analytical tracing of polarized light rays with accuracy comparable to numerical ray-tracing for Kerr black holes with spin parameters up to $a_{*}\simeq 0.94$, while being over four orders of magnitude faster Loktev, Vladislav et al. (2024).
@@
The primary goal of the present paper is to overcome this limitation by extending and generalizing our previous analytical framework to describe photon trajectories in generic static and spherically symmetric spacetimes. The resulting formulas provide a consistent description of direct rays—those completing less than half an orbit around the compact object—within a broad range of gravitational configurations, thereby broadening the applicability of semi-analytical modeling in strong-field regimes.
The organization of this paper is as follows. Section II presents the analytical approximations (23) and (24) together with their derivation. In Sec. III, we test the accuracy of these approximations. Section III.1 applies them to parameterized metrics such as the nonrotating Johannsen–Psaltis Johannsen and Psaltis (2011) and the Rezzolla–Zhidenko Rezzolla and Zhidenko (2014) metrics. Section III.2 considers a parameterization of the Einstein–Maxwell–dilaton–axion metric García et al. (1995). In Sec. IV, we present two applications of the approximations: Sec. IV.1 generates equatorial isoradial curves for the above metrics, while Sec. IV.2 computes the $QU$ loop diagrams produced by synchrotron-emitting hotspots in the Einstein–Maxwell–dilaton–axion spacetime. We conclude with some final remarks in Sec. V. Throughout this paper we adopt the (,+,+,+)(-,+,+,+) signature and geometrized units with $G=c=1$.
@@ II.1 Preliminaries
Figure 1: Geometry of a light ray traveling from the vicinity of a compact object to a distant observer. The ray reaches the observer, located in the asymptotic region ($\psi=0$), with an impact parameter $b$. The ray’s point of closest approach (periastron) occurs at the radial coordinate $r=p$. At a given coordinate $r=R$, $\alpha$ represents the angle between the radial direction and the light ray’s path.
@@ II.2 Approximation formulas for more general metrics
$ds^{2}=-A(r)dt^{2}+B(r)dr^{2}+r^{2}(d\theta^{2}+\sin^{2}\theta d\phi^{2}).$ (4)
Because the orbits of interest are confined to a plane, we can, without loss of generality, perform an angular coordinate transformation $(\theta,\phi)\rightarrow(\tilde{\theta},\psi)$ such that the light trajectory lies in the equatorial plane, defined by $\tilde{\theta}=\pi/2$. In this coordinate system, the metric takes the form
@@
As shown in Fig. 1, once the light ray passes the periastron pp—where the emission angle satisfies $\alpha=\pi/2$—and moves toward the asymptotic region, we have $\dot{\psi}<0$ and $\dot{r}>0$. In this region, the null geodesic equation can be written as
@@
Since the inverse relation is of interest—namely, expressing $x$ as a power series in yy—we consider an expansion of the form
@@
which for the particular case of a Schwarzschild metric reduces to the well-known Beloborodov formula Eq. (1). For this reason, in this work Eq. (24) will be referred to as the generalized Beloborodov formula. As will be shown below, this compact expression can also be successfully applied to describe light rays in a nontrivial class of astrophysical configurations as observed from infinity.
@@ III Validation of the Analytical Approximation
We now test the accuracy of Eqs. (23) and (24) by comparing them with the numerical results obtained from Eq. (7) for $\alpha<\pi/2$ and from Eq. (10) for $\alpha>\pi/2$. For this purpose, we consider two widely used parametric families describing black holes—the Johannsen–Psaltis and Rezzolla–Zhidenko metrics—together with a specific solution, the Einstein–Maxwell–dilaton–axion metric. The comparison is performed for a value of $R$ close to the event horizon $r_{H}$, specifically at $R=1.1\,r_{H}$, and for the set of radii $R=\{1.01\,r_{c},\,1.5\,r_{c},\,3\,r_{c},\,4.5\,r_{c}\}$, which correspond to different locations relative to the photon-sphere radius $r_{c}$, obtained as the solution of Chandrasekhar (1983); Cardoso et al. (2009):
@@
For all metrics considered, Approx. T denotes the approximation given by Eq. (23), corresponding to the Taylor expansion supplemented with an empirical term, whereas Approx. L refers to the linearized approximation derived from the generalized Beloborodov formula, Eq. (24). The results are compared with the numerical integration of $\psi$, from which the inverse relation $\alpha=\alpha(\psi)$ is obtained numerically, using Eq. (7) for $\alpha<\pi/2$ and Eq. (10) for $\alpha>\pi/2$.
@@ III.1 Metrics families
III.1.1 Johannsen-Psaltis family
@@
Figure 2: This figure compares the numerical integration of the exact relation $\cos\psi$ to $\cos\alpha$ (labeled ‘Exact’) with the analytical approximations from Eq. (23) (‘Approx. T’) and (24) (‘Approx. L’). The values $r_{H}$ and $r_{c}$ denote the radial coordinates of the event horizon and photon sphere, respectively, for the Johannsen-Psaltis metric. The bottom panel shows the relative error $\delta\alpha/\alpha\%$. Solid lines indicate the error for ‘Approx. T’ (vs ‘Exact’), while dashed lines indicate the error for ‘Approx. L’ (vs ‘Exact’).
@@
For the subsequent numerical calculations, terms up to the fourth order in $k$ in the function $h(r)$, as defined in Eq. (27), are retained. Since the metrics are required to be asymptotically flat and to satisfy the constraints of the parametrized post-Newtonian (PPN) formalism Will (2006), the first three parameters must be set to zero Johannsen and Psaltis (2011), i.e., $\epsilon_{i}=0$ for $i=0,1,2$. In that case, $M$ coincides with the Arnowitt–Deser–Misner (ADM) mass of the black hole. In what follows, the approximation is tested for different values of $\epsilon_{3}$ and $\epsilon_{4}$, with $M$ fixed to $1$. Figure 2 shows the relation between $\cos\psi$ and $\cos\alpha$ obtained from our analytical approximations, in direct comparison with the results of the numerical integration of (9) and (10), which must be inverted to express $\alpha$ as a function of $\psi$. The curves corresponding to Eqs. (23) and (24) represent the two analytical approximations under evaluation. The lower panel displays the relative percentage error in $\alpha$, illustrating the deviation of each analytical model from the numerical results. As shown in Fig. 2, Eq. (23) is virtually indistinguishable from the numerical result for all values of $R$ and parameter choices considered, maintaining a relative error below $2\%$ even in the vicinity of the event horizon. In contrast, Eq. (24) gradually loses accuracy as $\psi\rightarrow\pi$; however, for $R>1.5\,r_{c}$ it still provides a good approximation, with the relative error remaining below $2\%$ for $\cos\psi>0.25$.
@@ III.1.2 Rezzolla-Zhidenko family
Another metric with a high parameter dependence and widely used in the literature is the Rezzolla-Zhidenko Rezzolla and Zhidenko (2014) (RZ) metric, commonly used to test the strong field regime through black hole shadows Younsi et al. (2016b); Konoplya and Zhidenko (2019); Völkel et al. (2021); Mizuno et al. (2018), as well as parameterization of metrics for alternative theories of gravity De Laurentis et al. (2018); Kokkotas et al. (2017); Hennigar et al. (2018) and, more recently, the study of quasi-normal modes and modifications of the black hole radiation spectrum Dubinsky and Zinhailo (2025); Bolokhov and Skvortsova (2025); Siqueira et al. (2025). In contrast to similar approaches that rely on a Taylor expansion of the metric functions $A(r)$ and $B(r)$ in powers of $M/r$ (with $M$ the ADM mass), the authors employ a continued-fraction expansion of these functions. This representation offers improved convergence behavior and enables an accurate approximation of a nontrivial number of known metric solutions from various alternative theories of gravity with only a small set of coefficients. As shown by the authors, determining these coefficients from observations of near-horizon phenomena provides an effective way to constrain and compare different metric theories of gravity. These functions depend on an infinite set of dimensionless parameters: $\epsilon=(2M-r_{H})/r_{H}$, which measures deviations from the Schwarzschild radius; $a_{0},a_{1},a_{2},\ldots$, representing the corrections to $A(r)$; and $b_{0},b_{1},b_{2},\ldots$, representing the corrections to $B(r)$. As in the JP metric, both $a_{0}$ and $b_{0}$ are negligible within the PPN framework. Remarkably, Ref. Konoplya and Zhidenko (2020) showed that only a small set of parameters is needed to probe the strong-field regime. In particular, truncating the continued-fraction expansion of $A(r)$ and $B(r)$ at second order leaves $\epsilon$, $a_{1}$, $a_{2}$, $b_{1}$, and $b_{2}$ as free parameters, which suffice to describe astrophysical observables associated with near-horizon processes. As before, the ADM mass is fixed to unity. Under these considerations, the metric functions can be written as
@@
Figure 3 displays the relation between $\cos\psi$ and $\cos\alpha$ for five emission radii, comparing the numerical integration (solid) with the analytical approximations in Eqs. (23) (Approx. T, dotted) and (24) (Approx. L, dashed). The three panels correspond to distinct choices of the continued–fraction parameters $(\epsilon;a_{1},a_{2};b_{1},b_{2})$, with the associated $(r_{H},r_{c})$ indicated on top. In all cases, Approx. T tracks the numerical curves extremely well across the full range of $\cos\psi$, with a relative error that remains at the level of $\lesssim 2\%$ for all radii considered, including the near–horizon case $R=1.1\,r_{H}$ and the photon–sphere–proximate case $R=1.01\,r_{c}$. As expected, both approximations converge to the numerical result in the weak–deflection regime ($\cos\psi\!\to\!1$), where the curves nearly overlap.
The linearized scheme (Approx. L) behaves well in the moderate–bending regime, providing an accurate description for $\cos\psi>0$ and $R\gtrsim 1.5\,r_{c}$. Its accuracy degrades systematically as $\psi\to\pi$ (i.e., $\cos\psi\to-1$), particularly for the smallest radii ($R=1.1\,r_{H}$ and $R=1.01\,r_{c}$), where the dashed curves depart most visibly from the solid ones. The bottom panels quantify these trends: the error of Approx. T stays within a few percent and shows smooth, parameter–dependent sign changes, whereas Approx. L exhibits larger biases at large deflection but remains competitive for $\cos\psi\gtrsim 0$ and larger $R$. Importantly, the qualitative conclusions are robust across the different parameter sets: varying $(\epsilon;a_{1},a_{2};b_{1},b_{2})$ does not alter the observed hierarchy in performance between Approx. T and Approx. L.

### v2

@@ I Introduction
Among these approaches, the Beloborodov formula (Eq. (1) in Beloborodov (2002)) stands out. In the Schwarzschild spacetime, it provides a relation between the emission angle of a light ray; measured with respect to the radial direction; and the angular (and radial) position at which the same point is observed in the asymptotic region. Despite its success, this formula has some intrinsic limitations: it loses accuracy for emission angles exceeding $90^{\circ}$, it assumes exact spherical symmetry (thus neglecting spin-induced spacetime deformations), and it is restricted to the Schwarzschild metric.
Some of these shortcomings have been mitigated in later works. In particular, Poutanen Poutanen (2020) proposed a refined expression; still within the Schwarzschild geometry; that significantly improves the treatment of large-angle emission and has been successfully applied in various astrophysical contexts.
Moreover, comparative studies between approximate analytical treatments of photon propagation in a non-spinning Schwarzschild spacetime and full Kerr ray-tracing simulations indicate that, for moderate black-hole spins, both the resulting direct images and their associated polarization signatures remain in close quantitative agreement Cárdenas-Avendaño et al. (2023); Loktev, Vladislav et al. (2024). Notably, the Poutanen approximation has been implemented in the ARTPOL algorithm, which enables fast analytical tracing of polarized light rays with accuracy comparable to numerical ray-tracing for Kerr black holes with spin parameters up to $a_{*}\simeq 0.94$, while being over four orders of magnitude faster Loktev, Vladislav et al. (2024).
@@
The primary goal of this paper is to overcome this limitation by extending and generalizing our previous analytical framework to describe photon trajectories in generic static and spherically symmetric spacetimes. The resulting expressions provide a consistent treatment of direct rays; namely, those completing less than half an orbit around the compact object, for which the emission angle $\psi$ of a emission point $p$ as shown in Fig. 1 remains smaller than $\pi$. A more detailed discussion of this geometric configuration will be presented in the next section. Consequently, across a broad class of gravitational backgrounds, the new formulas significantly enhance the applicability of semi-analytical modeling in the strong-field regime.
The organization of this paper is as follows. Section II presents the analytical approximations (23) and (24) together with their derivation. In Sec. III, we test the accuracy of these approximations. Section III.1 applies them to parameterized metrics such as the nonrotating Johannsen–Psaltis Johannsen and Psaltis (2011) and the Rezzolla–Zhidenko Rezzolla and Zhidenko (2014) metrics. Section III.2 considers a parameterization of the Einstein–Maxwell–dilaton–axion metric García et al. (1995). In Sec. IV, we present two applications of the approximations: Subsec. IV.1 generates equatorial isoradial curves for the above metrics, while Subsec. IV.2 computes the $QU$ loop diagrams produced by synchrotron-emitting hotspots in the Einstein-Maxwell-dilaton-axion spacetime. We conclude with some final remarks in Sec. V. Throughout this paper we adopt the (,+,+,+)(-,+,+,+) signature and geometrized units with $G=c=1$.
@@ II.1 Preliminaries
Figure 1: Figure reproduced from Ref. Claros and Gallo (2024), Fig. 1. It illustrates a light ray traveling from the vicinity of a compact object to a distant observer. The ray reaches the observer, located in the asymptotic region ($\psi=0$), with an impact parameter $b$. The ray’s point of closest approach (periastron) occurs at the radial coordinate $r=p$. At a given coordinate $r=R$, $\alpha$ represents the angle between the radial direction and the light ray’s path.
@@ II.2 Approximation formulas for more general metrics
$ds^{2}=-A(r)dt^{2}+B(r)dr^{2}+r^{2}(d\theta^{2}+\sin^{2}\theta d\phi^{2});$ (4)
which differs from Eq. (3) by allowing for a more general function $B(r)$, instead of imposing $B(r)=A^{-1}(r)$. Because the orbits of interest are confined to a plane, we can, without loss of generality, perform an angular coordinate transformation $(\theta,\phi)\rightarrow(\tilde{\theta},\psi)$ such that the light trajectory lies in the equatorial plane, defined by $\tilde{\theta}=\pi/2$. In this coordinate system, the metric takes the form
@@
As shown in Fig. 1, once the light ray passes the periastron $p$-where the emission angle satisfies $\alpha=\pi/2$-and moves toward the asymptotic region, we have $\dot{\psi}<0$ and $\dot{r}>0$. In this region, the null geodesic equation can be written as
@@
Since the inverse relation is of interest-namely, expressing $x$ as a power series in $y$-we consider an expansion of the form
@@
which for the particular case of a Schwarzschild metric reduces to the well-known Beloborodov formula Eq. (1). For this reason, in this work Eq. (24) will be referred to as the generalized Beloborodov formula. As will be shown below, this compact expression can also be successfully applied to describe light rays in a nontrivial class of astrophysical configurations as observed from infinity. What is most interesting and surprising about all these analytical approximations (including the Beloborodov and Poutanen formulas) is that they provide an approximate relation between $\alpha$ and $\psi$ that remains valid regardless of whether the orbit passes through periastron. In numerical integrations, by contrast, one must first locate the periastron and determine the value $\psi_{\text{max}}$, so that the integral in Eq.(9) is used for $\psi\leq\psi_{\text{max}}$ and the expression in Eq.(10) is used for $\psi>\psi_{\text{max}}$. Remarkably, in the analytical approximations the relation between $\psi$ and $\alpha$ holds throughout the entire trajectory; as long as $\psi\lesssim\pi$; without the need to compute either the periaxis $p$ or $\psi_{\text{max}}$. For this reason, as we will see in the comparisons between numerical and analytical results in the next section, the numerical integration must be performed in two segments, whereas the analytical expression is single and given by Eqs.(23) or (24).
@@ III Validation of the Analytical Approximation
We now test the accuracy of Eqs. (23) and (24) by comparing them with the numerical results obtained from Eq. (7) for $\alpha<\pi/2$ and from Eq. (10) for $\alpha>\pi/2$. For this purpose, we consider two widely used parametric families describing black holes-the Johannsen–Psaltis and Rezzolla–Zhidenko metrics-together with a specific solution, the Einstein–Maxwell–dilaton–axion metric. The comparison is performed for a value of $R$ close to the event horizon $r_{H}$, specifically at $R=1.1\,r_{H}$, and for the set of radii $R=\{1.01\,r_{c},\,1.5\,r_{c},\,3\,r_{c},\,4.5\,r_{c}\}$, which correspond to different locations relative to the photon-sphere radius $r_{c}$, obtained as the solution of Chandrasekhar (1983); Cardoso et al. (2009):
@@
For all metrics considered, ‘Approx. T’ denotes the approximation given by Eq. (23), corresponding to the Taylor expansion supplemented with an empirical term, whereas ‘Approx. L’ refers to the linearized approximation derived from the generalized Beloborodov formula, Eq. (24). The results are compared with the numerical integration of $\psi$, from which the inverse relation $\alpha=\alpha(\psi)$ is obtained numerically, using Eq. (7) for $\alpha<\pi/2$ and Eq. (10) for $\alpha>\pi/2$.
@@ III.1 Metrics families
III.1.1 Non-spinning subclass of the Johannsen-Psaltis family
@@
Figure 2: This figure compares the numerical integration of the exact relation between $\cos\psi$ and $\cos\alpha$ (labeled ‘Exact’) with the analytical approximations given by Eqs. (23) (‘Approx. T’) and (24) (‘Approx. L’). As explained in the main text, the numerical integration is performed using the two expressions (9) and (10), as appropriate, and joined at the periaxis where $\cos\alpha=0$ and $\psi=\psi_{\text{max}}$. For the analytical approximation, Eqs. (23) or (24) are invoked directly over the entire range. The values $r_{H}$ and $r_{c}$ denote the radial coordinates of the event horizon and the photon sphere, respectively, for the Johannsen–Psaltis metric. The bottom panel shows the relative error $\delta\alpha/\alpha\,[\%]$. Solid lines indicate the error for ‘Approx. T’ (vs ‘Exact’), while dashed lines indicate the error for ‘Approx. L’ (vs ‘Exact’).
@@
For the subsequent numerical calculations, terms up to the fourth order in $k$ in the function $h(r)$, as defined in Eq. (27), are retained. Since the metrics are required to be asymptotically flat and to satisfy the constraints of the parametrized post-Newtonian (PPN) formalism Will (2006), the first three parameters must be set to zero Johannsen and Psaltis (2011), i.e., $\epsilon_{i}=0$ for $i=0,1,2$. In that case, $M$ coincides with the Arnowitt–Deser–Misner (ADM) mass of the black hole. In what follows, the approximation is tested for different values of $\epsilon_{3}$ and $\epsilon_{4}$, with $M$ fixed to $1$. Figure 2 shows the relation between $\cos\psi$ and $\cos\alpha$ obtained from our analytical approximations, in direct comparison with the results of the numerical integration of (9) and (10), which must be inverted to express $\alpha$ as a function of $\psi$. The curves corresponding to Eqs. (23) and (24) represent the two analytical approximations under evaluation. The lower panel displays the relative percentage error in $\alpha$, illustrating the deviation of each analytical model from the numerical results. As shown in Fig. 2, Eq. (23) is virtually indistinguishable from the numerical result for all values of $R$ and parameter choices considered, maintaining a relative error below $3\%$ even in the vicinity of the event horizon. In contrast, Eq. (24) gradually loses accuracy as $\psi\rightarrow\pi$; however, for $R>1.5\,r_{c}$ it still provides a good approximation, with the relative error remaining below $2\%$ for $\cos\psi>0.25$.
@@ III.1.2 Rezzolla-Zhidenko family
Another metric with a high parameter dependence and widely used in the literature is the Rezzolla-Zhidenko Rezzolla and Zhidenko (2014) (RZ) metric, commonly used to test the strong field regime through black hole shadows and strong gravitational lensingYounsi et al. (2016b); Konoplya and Zhidenko (2019); Völkel et al. (2021); Mizuno et al. (2018); Feleppa et al. (2025), as well as parameterization of metrics for alternative theories of gravity De Laurentis et al. (2018); Kokkotas et al. (2017); Hennigar et al. (2018) and, more recently, the study of quasi-normal modes and modifications of the black hole radiation spectrum Dubinsky and Zinhailo (2025); Bolokhov and Skvortsova (2025); Siqueira et al. (2025). In contrast to similar approaches that rely on a Taylor expansion of the metric functions $A(r)$ and $B(r)$ in powers of $M/r$ (with $M$ the ADM mass), the authors employ a continued-fraction expansion of these functions. This representation offers improved convergence behavior and enables an accurate approximation of a nontrivial number of known metric solutions from various alternative theories of gravity with only a small set of coefficients. As shown by the authors, determining these coefficients from observations of near-horizon phenomena provides an effective way to constrain and compare different metric theories of gravity. These functions depend on an infinite set of dimensionless parameters: $\epsilon=(2M-r_{H})/r_{H}$, which measures deviations from the Schwarzschild radius; $a_{0},a_{1},a_{2},\ldots$, which represent corrections to the Schwarzschild expression for $A(r)$; and $b_{0},b_{1},b_{2},\ldots$, which represent corrections to the Schwarzschild expression for $B(r)$. As in the JP metric, both $a_{0}$ and $b_{0}$ are negligible within the PPN framework. Remarkably, Ref. Konoplya and Zhidenko (2020) showed that only a small set of parameters is needed to probe the strong-field regime. In particular, truncating the continued-fraction expansion of $A(r)$ and $B(r)$ at second order leaves $\epsilon$, $a_{1}$, $a_{2}$, $b_{1}$, and $b_{2}$ as free parameters, which suffice to describe astrophysical observables associated with near-horizon processes. As before, the ADM mass is fixed to unity. Under these considerations, the metric functions can be written as
@@
Figure 3 displays the relation between $\cos\psi$ and $\cos\alpha$ for five emission radii, comparing the numerical integration (solid) with the analytical approximations in Eqs. (23) (‘Approx. T’, dotted) and (24) (‘Approx. L’, dashed). The three panels correspond to distinct choices of the continued–fraction parameters $(\epsilon;a_{1},a_{2};b_{1},b_{2})$, with the associated $(r_{H},r_{c})$ indicated on top. In all cases, ‘Approx. T’ tracks the numerical curves extremely well across the full range of $\cos\psi$, with a relative error that remains at the level of $\lesssim 2\%$ for all radii considered, including the near–horizon case $R=1.1\,r_{H}$ and the photon–sphere–proximate case $R=1.01\,r_{c}$. As expected, both approximations converge to the numerical result in the weak–deflection regime ($\cos\psi\!\to\!1$), where the curves nearly overlap.
The linearized scheme (‘Approx. L’) behaves well in the moderate–bending regime, providing an accurate description for $\cos\psi>0$ and $R\gtrsim 1.5\,r_{c}$. Its accuracy degrades systematically as $\psi\to\pi$ (i.e., $\cos\psi\to-1$), particularly for the smallest radii ($R=1.1\,r_{H}$ and $R=1.01\,r_{c}$), where the dashed curves depart most visibly from the solid ones. The bottom panels quantify these trends: the error of ‘Approx. T’ stays within a few percent and shows smooth, parameter–dependent sign changes, whereas ‘Approx. L’ exhibits larger biases at large deflection but remains competitive for $\cos\psi\gtrsim 0$ and larger $R$. Importantly, the qualitative conclusions are robust across the different parameter sets: varying $(\epsilon;a_{1},a_{2};b_{1},b_{2})$ does not alter the observed hierarchy in performance between ‘Approx. T’ and ‘Approx. L’.